: Consider a sphere of radius 5 mm placed in free space containing a total amount of charge ‘Q’. The electric potential due to the distribution of the charge is V = 7 R2 Sin θ Cos ϕ. Find out the numeric value of the charge Q enclosed by the surface.
1
Expert's answer
2020-05-25T10:49:01-0400

Let the potential be:


"V = 7r^2 \sin\theta\cos \varphi"


According to the Maxwell's equations, the charge density will be:


"\bigtriangleup V = -\dfrac{\rho}{\varepsilon_0}"

where ρ is the charge density and ε0 ​ is the dielectric constant.

The Laplace operator in the spherical coordinates will have the following view:


"\Delta V = \frac{1}{r^2} \frac{\partial}{\partial r} \left(r^2 \frac{\partial V}{\partial r} \right) + \frac{1}{r^2 \sin \theta} \frac{\partial}{\partial \theta} \left(\sin \theta \frac{\partial V}{\partial \theta} \right) + \frac{1}{r^2 \sin^2 \theta} \frac{\partial^2 V}{\partial \varphi^2}"

The first term:


"\frac{1}{r^2} \frac{\partial}{\partial r} \left(r^2 \frac{\partial V}{\partial r} \right) = \frac{1}{r^2} \frac{\partial}{\partial r} \left(r^2 \cdot14r\sin\theta\cos \varphi \right) = \\
=\frac{1}{r^2}\cdot42r^2\sin\theta\cos \varphi =42\sin\theta\cos \varphi"

The second term:


"\frac{1}{r^2 \sin \theta} \frac{\partial}{\partial \theta} \left(\sin \theta \frac{\partial V}{\partial \theta} \right) =\frac{1}{r^2 \sin \theta} \frac{\partial}{\partial \theta} \left(\sin \theta \cdot 7r^2 \cos\theta\cos \varphi \right) =\\
=\frac{7\cos \varphi}{2 \sin \theta} \frac{\partial}{\partial \theta} \left(\sin2\theta \right) = 
\frac{7\cos \varphi\cos\theta}{\sin \theta} =7\cos\varphi\tan\theta"


The third term:


"\frac{1}{r^2 \sin^2 \theta} \frac{\partial^2 V}{\partial \varphi^2} = \frac{7r^2\sin\theta}{r^2 \sin^2 \theta} \frac{\partial^2 \cos\varphi}{\partial \varphi^2} =\frac{7}{\sin \theta} \frac{\partial^2 \cos\varphi}{\partial \varphi^2} = -\frac{7\cos\varphi}{\sin \theta}"

Putting it all together:


"\rho = 7\varepsilon_0 \left( \frac{\cos\varphi}{\sin \theta}- 6\sin\theta\cos \varphi - \cos\varphi\tan\theta \right)"

Now to find the total charge, one should integrate this density over whole sphere:


"Q = \iiint\rho dV = \iiint\rho r^2 \sin\theta drd\theta d\varphi"

The integration w.r. to r will give the multiplier "r_0^3\/3" , where r0 = 5 mm is the radius of the sphere, because the charge density does not depend on the radial coordinate.

Integrating the first term of density:


"\iint\frac{\cos\varphi}{\sin \theta} \sin\theta d\theta d\varphi = 4\int^\pi _0d\theta\int^{\pi/2}_0 \cos\varphi d\varphi = 
4\pi\sin\varphi|^{\pi/2}_0 = 4\pi"

Integrating the second term of density:


"-\iint6\sin^2\theta\cos \varphi d\theta d\varphi =-6 \cdot4\int^\pi _0 \sin^2\theta d\theta \int^{\pi/2}_0 \cos\varphi d\varphi = -6\cdot\dfrac{\pi}{2}\cdot 4\pi = -48\pi^2"

Integrating the third term of density:


"-\iint \tan\theta\cos \varphi \sin\theta d\theta d\varphi =- \int^\pi _0 \tan\theta\sin\theta d\theta \int^{2\pi}_0 \cos\varphi d\varphi = 0"

Finally:


"Q = 7\varepsilon_0 \dfrac{r_0^3}{3} \left(4\pi - 48\pi^2 \right) = 7\cdot8.85\cdot 10^{-12} \dfrac{0.005^3}{3} \left(4\pi - 48\pi^2 \right) = -1.2\cdot 10^{-15} C"

Answer. Q = -1.2*10^(-15) C